OpenStudy (angiepangie0726):

Fan and Medal :) Solve by factoring x3 − x2 − 24x − 36 = 0.

OpenStudy (angiepangie0726):

@sshayer

OpenStudy (sshayer):

put x=-2 -8-4+48-36=48-48=0 one solution is x=2

OpenStudy (angiepangie0726):

The others are 3 and -6 right?? :)

OpenStudy (sshayer):

correction one solution is x=-2 divide by 2 

OpenStudy (sshayer):

you are correct.